OpenStudy (anonymous):

This stuff is a foreign language to me can someone help me with, Write the slope-intercept equation for the line that passes through (4, -9) and has slope -3/4.

OpenStudy (saifoo.khan):

Satellite will help Directly from the TOP.

OpenStudy (anonymous):

hola come estas

OpenStudy (anonymous):

use \[y-y_1=m(x-x_1)\] with \[m=-\frac{3}{4}, x_1=4,y_1=-9\]

OpenStudy (anonymous):

put them in place and get \[y+9=-\frac{3}{4}(x-4)\]

OpenStudy (anonymous):

multiply out and get \[y+9=-\frac{3}{4}x+3\] subtract 9 to get \[y=-\frac{3}{4}x-6\]

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] is the "point-slope" form \[y=mx+b\] is the "slope - intercept " form